OpenStudy (anonymous):

Find the intecepts and use them to graph the equation 2x+6y=12

OpenStudy (anonymous):

make y the subect of the formula eg. y=mX+c then the whole number will be your y-intercept (C) to find x-intercept let y=o

OpenStudy (anonymous):

I'm sorry still confused @ Physicboffin

OpenStudy (anonymous):

6y=-2X+12 Y= -2/6 X +2 therefore your y-intercept is 2 (0,2) then x -intercept let y=o o=-2/6 X+2 -2=-2/6X X=6 therefore x-intercept (6,0)

OpenStudy (anonymous):

Oh now I get it thanks, your the best of the class this year lmbo!!!

OpenStudy (anonymous):

Bazinga

OpenStudy (anonymous):

I guess.... Bazinga lol